OpenStudy (anonymous):

what is 7*10^3 over 2*10^7

OpenStudy (anonymous):

0.00035

OpenStudy (anonymous):

well its multiple choice and thats not one of the answers thanks anyway

OpenStudy (anonymous):

A. 3.5*10^4 B.3.5*10^-4 C.3.5*10^10 D.3.5*10^-5

OpenStudy (unklerhaukus):

\[\frac{7\times 10^3}{2\times 10^7}\]\[=\frac{7}{2}\times 10^{-4}\]\[=3.5\times 10^{-4}\]\[=0.00035\]

OpenStudy (anonymous):

my choices are listen above im not sure which one it is...

OpenStudy (anonymous):

b

OpenStudy (unklerhaukus):

B

OpenStudy (anonymous):

thanks so much